Four marbles are drawn from a bag containing six green and five purple marbles.
What is the probability that all four marbles are purple?
i) if the marble is replaced after each draw?
ii) if the marbles are not replaced?

Answer by Boreal(15235)

if replaced, it is (5/11)^4=0.04269
if not replaced, it is (5/11)(4/10)(3/9)(2/8)=120/7920=0.01515 (repeat)
also
11C4 are the number of outcomes, or 330
5C4 are the ways to choose 4 purple marbles =5
it is 5/330 or 1/66 in this scenario. That is 0.01515...